Circular No. 7877

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002cf IN NGC 4786 B. Beutler and W. D. Li, University of California, report the LOTOSS

(cf. IAUC 7514) discovery of an apparent supernova on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Apr. 13.3 (mag ∼ 17.0) and 14.3 UT (mag ∼ 16.9). SN 2002cf is located at α = 12^{h}54^{m}31^{s}.30, δ = −06^{o}51′24′′.8 (equinox 2000.0), which is 16′′.4 west and 9′′.4 north of the nucleus of NGC 4786. A KAIT image taken on Mar. 12.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002cg IN UGC 10415 Li also reports the LOTOSS discovery of an apparent supernova in

unfiltered KAIT images taken on Apr. 13.5 (mag ∼ 17.8) and 14.4 UT (mag ∼ 17.4). SN 2002cg is located at α = 16^{h}29^{m}00^{s}.48, δ = +41^{o}17′00′′.4 (equinox 2000.0), which is 1′′.2 east and 2′′.9 south of the nucleus of UGC 10415. KAIT images taken on Mar. 28.5 (limiting mag ∼ 18.0) and 2001 Aug 16.3 (limiting mag ∼ 19.5) showed nothing at this position.

SUPERNOVA 2002cc IN ANONYMOUS GALAXY T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002cc (cf. IAUC 7871), obtained by M. Calkins on Apr. 10.19 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova about a week past maximum. Narrow Hα emission from a separate exposure of the nucleus indicates a recession velocity of 19 950 km/s for the host galaxy. Using this value, the supernova expansion velocity is ∼ 10 200 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (cf. IAUC 7774) was 9 ± 2 days after maximum light.

XTE J1908+094 P. Garnavich and J. Quinn, University of Notre Dame; and P. Callanan,

University College, Cork, report that images (exposures 1500 s in I, 2100 s in R), obtained with the 1.8-m Vatican Advanced Technology Telescope on Apr. 8 and 9 UT, show no obvious new sources and no strongly varying sources within the stated position error circles of this transient (IAUC 7856, 7873) when compared with the red Digitized Sky Survey. No sources (limiting mag I = 22.0; astrometric accuracy 0′′.2 via six USNO-A2.0 stars) are found within 3′′ of the radio position (IAUC 7874).

2002 April 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green